# Towards a Reward-Free Reinforcement Learning Framework for Vehicle Control

### v1

@@ III-B Reward-Free State-Guided Policy Network (RFSGPN)
Next, we introduce the design and implementation of the RFSGPN. The RFSGPN consists of multiple convolutional layers, batch normalization layers, and fully connected layers. The main role of the convolutional layers is to extract useful features from the input states, which is crucial for understanding the environment and making decisions. To more effectively process features, we design a feature extraction network and obtain the extracted features:
@@
where $s_{t}$ is the current input state, initialized as the provided initial state, and subsequent states are obtained as the next state after interaction with the environment. $Q(\cdot)$ is a feature extraction network composed of three convolutional layers. After each convolutional layer, there is a batch normalization layer to improve the learning ability and stability of the network [50]. When the input of RFSGPN is numerical data, it is not suitable to use convolutional layers to extract features. we change the convolution layers in (6) to linear layers. Subsequently, the fully connected layer transforms the features extracted by the convolutional layers into a mapping of the action space:
$\begin{gathered}h_{t}=\text{ReLU}(W_{1}f_{3}+b_{1}),\\ q_{t}=\text{ReLU}(W_{2}h_{t}+b_{2}),\end{gathered}$ (7)
where $W_{1}$ and $W_{2}$ are the weight matrices of the fully connected layer, $b_{1}$ and $b_{2}$ are the bias terms, $q_{t}$ is the output action probability distribution. In order to ensure that the output of the network matches the size of the action space, we set the number of output nodes equal to the size of the action space. Finally, the action probability distribution of the network output can be expressed as
$\pi(a_{t}\mid s_{t})=\mathrm{softmax}(q_{t}).$ (8)
In this way, the network can generate a probability distribution for action selection based on the extracted features.
To optimize the policy network in the absence of reward signals, we adopt a new approach. In this situation without manually designed reward signals, selecting the action with the highest probability can make the policy more stable and reduce fluctuations caused by uncertainty. For the continuous action space, we discretize it and put the discretized actions into the discrete action pool. Then, we obtain the index $i_{t}$ corresponding to the maximum action by the following method, and then obtain the corresponding action $a_{t}$ from the discrete action pool:
$\begin{gathered}i_{t}=\mathop{\arg\max}\pi(a_{t}\mid s_{t}),\\ a_{t}=\text{ActionPool}[i_{t}],\end{gathered}$ (9)
where ActionPool is discrete action pool. This action $a_{t}$ serves two purposes. First, the action is used to interact with the real environment to obtain the next state $s_{t+1}$. At next step $t+1$, this new state replaces $s_{t}$ in (6). The initial state of each learning step uses the real state output by the environment, rather than the predicted state, which helps correct errors in the prediction model during policy learning. Second, the action is injected into the TSPN with frozen parameters to predict the next state:
$s_{t+1}^{pre}=\text{TargetNet}_{\text{frozen}}(s_{t},a_{t}),$ (10)
where $\text{TargetNet}_{\text{frozen}}$ represents the TSPN with frozen parameters. When training the RFSGPN, direct interaction with the environment will not be able to effectively backpropagate without the TSPN. Instead of using manually designed reward signal for optimizing the policy network, we use the difference between the next state from expert data $s^{exp}_{t+1}$ and the predicted next state $s^{pre}_{t+1}$ as our loss. We employ the following mean squared error loss function:
@@
where $N$ is the batch size. In this way, our network training process leverage the advantage of TSPN in state prediction accuracy and advantage of RFSGPN in learning policy, achieving effective learning without manually designing reward signals. Hence, our method is application to environments where reward signals are difficult to design or missing.
Additionally, for the RFSGPN of numerical data, it is not suitable to use convolutional layers to extract features, so in terms of structure, we change the original convolutional layers to linear layers.
@@ IV Experiments and Discussion
  • How do different parameter settings affect the performance of the algorithm, and can it maintain its stability under different training conditions?
@@ IV-A1 Data Collection
The data collection is executed by a data collector equipped with an initialized policy network. It engages in random interactions with the environment, collecting observational data and predicting actions through an untrained, initialized policy agent. The purpose of this process is to gather a series of state-action-next state triplets, which will be used for the training of TSPN.
During the data collection process, we use a temperature parameter to adjust the randomness of the actions selected by the agent. We set the temperature parameter to 1.0 to maintain a balance between exploration and exploitation. The setting of this parameter is an important experimental factor, which we study in detail in the subsequent section IV-D.
@@ IV-A2 Data processing
IV-A3 Network Training
In our framework, training is divided into two stages: independent training of the TSPN and subsequent training of the RFSGPN, with the latter being performed on the basis of fixing the parameters of the former.
During the training of the TSPN, we use Scheduled Sampling techniques and the Adafactor optimization algorithm. Scheduled Sampling is employed to produce the input to the network by determining, through a probability $\epsilon$ , whether to use the actual state sequence or the network’s own predictive output. We initially set $\epsilon$ to a value close to 1 and gradually begin to decay it within the first two iterations; in the subsequent iterations, we set $\epsilon$ to 0.5. The purpose of this setting is to allow the network’s input to be the real state sequence at the beginning of training, but as training progresses, the network’s input gradually transitions from the real state sequence to the network’s own prediction results, which helps to improve the network’s stability. We use mean squared error loss function, training the network by optimizing this loss function.
During the training of the RFSGPN, we employ the Adafactor optimization algorithm along with the mean square error loss function for parameter update. Specifically, with the TSPN parameters fixed, the agent starts from the initial state, predicts the action and interacts with the environment to obtain the next state, and sends the action to the TSPN to obtain the next predicted state. The mean square error loss function is used to calculate the error between the predicted state and the expert state, and the parameters of the RFSGPN are updated by the Adafactor algorithm.
@@ IV-B Baselines
Dream: A meta reinforcement learning framework proposed by Liu et al. [23]. This framework avoids the problem of local optimum by optimizing exploration and exploitation separately, allowing the agent to learn quickly on new tasks. Notably, it does not depend on explicit reward signals, thereby regarded as a reward-free reinforcement learning approach.
IV-C Comprehensive Environmental Experiment Analysis
@@ IV-C1 Carla Environment Experiment Analysis
$A=\{\delta,\tau\},$ (13)
here, we limit the vehicle’s steering angle $\delta$ and throttle $\tau$ to the range of [-0.8, 0.8] and [0.6, 1], respectively.
@@
Among them, $k_{1}$ and $k_{2}$ take the values of 0.5 and 0.1, respectively. Although we have defined rewards, they are only used to verify the effectiveness of our method in the final testing phase and are not utilized during the training process. Additionally, when the vehicle collides with an obstacle, reaches the destination, or is more than 15 meters away from the track, ’done’ becomes true, ending the current episode.
Evaluation: Here, we assess and compare the performance of our method with four baseline methods. We evaluate the performance of the agent under each method for 200 episodes and present the testing performance of different methods on this task in Table I. Correspondingly, Fig. 3 shows the test results for each episode during the testing period. In Table I, in addition to the maximum and average values, we also use the Inter-Quartile Mean (IQM) and the normalized IQM as evaluation metrics. These two indicators can effectively reduce the impact of extreme values on the average, thereby more accurately reflecting the central tendency of the test results. Table I and Fig. 3 indicate that, overall, RFRLF (Ours) has surpassed existing baselines in terms of maximum, average, and IQM metrics.
@@
From the results, it can be seen that the Import, Varibad and Dream algorithms do not achieve ideal performance. Although the Import algorithm reduces the dependence on the reward signal to a certain extent by utilizing the information of the task description, it is difficult to find an effective strategy based on this alone when the reward signal is completely missing. The Varibad algorithm involves a complex VAE model and meta-learning process, which may not be able to adapt to the reward-free environment and the autonomous driving scenario that requires rapid response, resulting in decision delays and failure to learn effective strategies. In the absence of such a reward signal, the Dream algorithm may require additional information or policies for effective exploration and learning. In addition, although IPL can simultaneously learn implicit rewards and optimal policies that meet the expert’s preferences, and can improve the learning process by collecting feedback data, it has a relatively good effect in this task, but expert preferences still introduce human errors. Compared with these methods, the RFRLF framework we proposed shows significant advantages. RFRLF optimizes the behavior of the agent and guides it to learn an effective policy without manually designing reward signals by combining the TSPN and the RFSGPN.
@@ IV-C2 Autocar Environment Experiment Analysis
Rewards: If the vehicle is moving forward, the reward is the current speed; if the vehicle is turning, the reward is half of the current speed (to prevent excessive turning); if the vehicle is making a half-circle turn, the reward is -45; if the vehicle goes off the track or has a collision, the reward is -150, and the episode ends. Like the previous experiment, although we have defined rewards, they are only used to verify the effectiveness of our method in the final testing phase and are not used in the training process.
@@
IV-D Performance Evaluation of Target State Prediction Network
@@
The key reason why the TSPN can achieve high-precision predictions lies in the action injection layer in its design. By injecting action information into the state features, the network is able to capture the impact of actions on state changes. In addition, the upsampling and downsampling modules not only effectively preserve the spatial resolution of the features, but also enable the network to capture richer contextual information, thereby further improving the prediction accuracy.
@@
IV-E Parameter Sensitivity Analysis
In the Carla environment, we implement 5 comparison schemes by setting various hyperparameters and making changes to components. In Table II, “Partial Freeze (PF)” indicates freezing some layers (input embedding layer and downsampling layer) of the TSPN. This is to analyze the impact of free adjustment of other layers on performance while retaining some feature extraction capabilities. In order to analyze the impact of temperature parameter “$T$=1.0” and scheduled sampling rate “$\epsilon$=0.5” on the RFRLF method, we also set $T$ in {0.5,2.0} and $\epsilon$ in {0.3,0.7} , where $T$ is the temperature parameter during data collection and $\epsilon$ is the scheduled sampling rate after the TSPN training epoch is greater than 2. These parameters are selected to observe the impact of smaller and larger temperatures and sampling rates on model exploration and training stability.
@@ IV-E Parameter Sensitivity Analysis
Overall, the performance of RFRLF is comparable to the baseline with altered hyperparameters and outperforms the baseline with partial component structure changes in terms of Mean Value and IQM. Specifically, by comparing with the “PF” baseline, it is found that after freezing part of the TSPN, the network performance does not change much. This may be because the frozen input embedding layer and downsampling layer effectively retain the network’s ability to extract input features, while allowing other unfrozen layers to still be adjusted and optimized during training, thereby ensuring the stability of the overall performance.
In the temperature parameter analysis, the “$T$=2.0” baseline performs best. This is because when training the TSPN, a higher temperature coefficient enhances exploration, enabling the agent to explore more states, thereby providing the TSPN with richer data. However, as the temperature coefficient increases, instability also increases, which means that the $T$ should not be too high, and a balance needs to be found between exploration and stability. Therefore, “$T$=1.0” as the default value is a suitable compromise. In addition, by comparing the “$T$=0.5”, “$T$=2.0”, “$\epsilon$=0.3”, and “$\epsilon$=0.7” baselines, we find that hyperparameters have a certain impact on the performance of RFRLF, but in general, RFRLF is not very sensitive to changes in hyperparameters. This shows that RFRLF has strong robustness within a certain range and can adapt to different hyperparameter settings while maintaining good performance.
@@ V Conclusion
This paper presents a novel Reward-Free Reinforcement Learning Framework (RFRLF), aimed at optimizing the agent’s policy without manually designed reward signals and without explicit action information. RFRLF consists of two main modules: the Target State Prediction Network (TSPN) and the Reward-Free State-Guided Policy Network (RFSGPN). The TSPN predicts the target state based on the agent’s current state and the actions it performs. The RFSGPN uses the output information of the TSPN and expert demonstration data that does not contain action and reward information to optimize the agent’s policy.
Experimental results show that RFRLF demonstrates superior performance compared to various reward-based and reward-free reinforcement learning methods in the Carla and Autocar environments. Our framework has surpassed baseline methods in terms of maximum value, average value, and IQM metrics, proving its effectiveness for vehicle control in reward-free environments.
Although RFRLF performs well in simulation environments, its performance still needs to be optimized in the real world to cope with more uncertainties. In the future, the application and development of RFRLF in real environments will be further promoted by introducing dynamic environmental factors and more efficient optimization algorithms.

### v2

@@ III-B Reward-Free State-Guided Policy Network (RFSGPN)
Next, we introduce the design and implementation of the RFSGPN. The RFSGPN extracts state features through a network $Q(\cdot)$ composed of convolutional layers, batch normalization, and fully connected layers, which is formulated by
@@
where $s_{t}$ is the current input state, initialized as the provided initial state, and subsequent states are obtained as the next state after interaction with the environment.[45]. When the input of RFSGPN is numerical data, it is not suitable to use convolutional layers to extract features. we change the convolution layers in (6) to fully-connected layers. Subsequently, a fully connected layer with the Gumbel-Softmax function generates the action probability distribution. For most reinforcement learning algorithms, since the environment dynamics are unknown, the discrepancy between the environment-generated states and target states cannot directly guide policy learning through backpropagation. However, by leveraging our pre-trained prediction network, states can propagate gradients through this network to optimize the policy. We therefore introduce a novel mechanism: directly feeding the policy distribution $\pi(a_{t}\mid s_{t})$ into the parameter-frozen Target State Prediction Network (TSPN) to predict the next state $s_{t+1}^{pre}$, then using the discrepancy between predicted and target states for policy optimization via backpropagation. This approach eliminates both the need for explicit action sampling and the requirement for value networks, fundamentally circumventing dependency on reward engineering. The process is formally expressed as:
$s_{t+1}^{pre}=\text{TargetNet}_{\text{frozen}}(s_{t},\pi(a_{t}\mid s_{t})),$ (7)
where $\text{TargetNet}_{\text{frozen}}$ represents the input embedding and feature extraction layers of the TSPN with frozen parameters, and $s_{t}$ represents the environment-generated state at each timestep. Here we use actual environment states $s_{t}$ rather than predicted states, which helps correct errors in the prediction model during policy learning. To obtain $s_{t}$ at each step, actions $a_{t}$ are sampled from $\pi(a_{t}\mid s_{t})$ and this process does not participate in backpropagation. For discrete action spaces, direct sampling is employed, whereas continuous action spaces require discretization prior to sampling. These implementation specifics are omitted here as they are non-essential to the core contributions of this work. After obtaining $s_{t+1}^{pre}$, we compute the following loss function to train the policy network:
@@
where $N$ is the batch size, $s_{t+1}^{exp}$ is the next state of the expert data. In this way, our network training process leverage the advantage of TSPN in state prediction and advantage of RFSGPN in policy learning, achieving effective learning without manually designing reward signals. Hence, our method is applicable to environments where reward signals are difficult to design or missing.
@@ IV Experiments and Discussion
  • How do different parameter settings affect the performance of the algorithm, and can it maintain its effectiveness under different training conditions?
@@ IV-A1 Data Collection
The data collection is executed by a data collector equipped with an initialized policy network. It engages in random interactions with the environment, collecting observational data and predicting actions through an untrained, initialized policy agent. The purpose of this process is to gather a series of (state, action, next state) triplets, which will be used for the training of TSPN.
During the data collection process, we use a temperature parameter to adjust the randomness of the actions selected by the agent. We set the temperature parameter to 1.0 to maintain a balance between exploration and exploitation. The setting of this parameter is an important experimental factor, which we study in detail in the subsequent section IV-F.
@@ IV-A2 Data processing
IV-A3 Network Training Configuration
In our framework, training is divided into two phases: independent training of the TSPN followed by the RFSGPN training, with the latter conducted while keeping the parameters of the former fixed. This section details the training parameter settings and specific implementations for both phases.
For the TSPN, we employ the Adafactor optimization algorithm with an initial learning rate of 0.001 using a cosine annealing strategy, a batch size of 64, and train for 100 epochs, each containing 1000 iterations. We utilize mean squared error as the loss function and implement Scheduled Sampling with an initial $\epsilon$ value of 0.9, which linearly decays to 0.5 during the first two training cycles and remains constant thereafter. This configuration enables the network to primarily rely on actual state sequences during the initial training phase, gradually transitioning to utilizing its own prediction results, effectively enhancing model stability.
For the RFSGPN training, we also utilize the Adafactor optimization algorithm but with a fixed learning rate of 0.0005, a batch size of 32, and train for 50 epochs with an early stopping strategy based on validation set performance. During the training process, we fix the parameters of the pre-trained TSPN and only update the policy network parameters, performing gradient backpropagation by computing the mean squared error between predicted states and expert states. The entire training process is implemented using the PyTorch 1.10.0 framework on an NVIDIA RTX 3090 GPU, with gradient clipping (maximum gradient norm of 1.0) to ensure training stability.
@@ IV-B Baselines
Dream: A meta reinforcement learning framework proposed by Liu et al. [23]. This framework avoids the problem of local optimum by optimizing exploration and exploitation separately, allowing the agent to learn quickly on new tasks. Notably, it does not depend on explicit reward signals, thereby selected as a baseline of our method.
IV-C Simulation Environmental Experiment Analysis
@@ IV-C1 Carla Environment Experiment Analysis
$A=\{\delta,\tau\}.$ (10)
Here we limit the vehicle’s steering angle $\delta$ and throttle $\tau$ to the range of [-0.8, 0.8] and [0.6, 1], respectively.
@@
Among them, $k_{1}$ and $k_{2}$ take the values of 0.5 and 0.1, respectively. Although we define rewards, they are only used to verify the effectiveness of our method in the final testing phase and are not utilized during the training process. Additionally, when the vehicle collides with an obstacle, reaches the destination, or is more than 15 meters away from the track, ’done’ becomes true, ending the current episode.
Evaluation: Here, we compare the performance of our method with four baseline methods. We evaluate the performance of the agent under each method for 200 episodes and present the testing performance of different methods on this task in Table I. Correspondingly, Fig. 3 shows the test results for each episode during the testing period. In Table I, in addition to the maximum and average values, we also use the Inter-Quartile Mean (IQM) and the normalized IQM as evaluation metrics. These two indicators can effectively reduce the impact of extreme values on the average, thereby more accurately reflecting the central tendency of the test results. Table I and Fig. 3 indicate that, overall, RFRLF (Ours) has surpassed existing baselines in terms of maximum, average, and IQM metrics.
@@
From the results, it can be seen that the Import, Varibad and Dream algorithms do not achieve ideal performance. Although the Import algorithm reduces the dependence on the reward signal to a certain extent by utilizing the information of the task description, it is difficult to find an effective strategy based on this alone when the reward signal is completely missing. The Varibad algorithm involves a complex VAE model and meta-learning process, which may not be able to adapt to the reward-free environment and the autonomous driving scenario that requires rapid response, resulting in decision delays and failure to learn effective strategies. In the absence of such a reward signal, the Dream algorithm may require additional information or policies for effective exploration and learning. In addition, although IPL can simultaneously learn implicit rewards and optimal policies that meet the expert’s preferences, and can improve the learning process by collecting feedback data, it has a relatively good effect in this task, but expert preference data still introduce human errors. Compared with these methods, the RFRLF framework we proposed shows significant advantages. RFRLF optimizes the behavior of the agent and guides it to learn an effective policy without manually designing reward signals by combining the TSPN and the RFSGPN.
@@ IV-C2 Autocar Environment Experiment Analysis
Rewards: If the vehicle is moving forward, the reward is the current speed; if the vehicle is turning, the reward is half of the current speed (to prevent excessive turning); if the vehicle is making a half-circle turn, the reward is -45; if the vehicle goes off the track or has a collision, the reward is -150, and the episode ends. Like the previous experiment, although we define rewards, they are only used to verify the effectiveness of our method in the final testing phase and are not used in the training process.
@@
IV-D Real-world Scenario Experimental Validation
To verify the effectiveness of the proposed RFRLF framework in real-world environments, this section constructs an actual experimental platform based on the TurboPi intelligent vision vehicle and conducts tests on an autonomous driving track. The experimental objective is to validate whether RFRLF can achieve stable control solely through expert state information without reward signals and expert action guidance.
Hardware Platform: The experiment employs the TurboPi intelligent vision vehicle, with core components including a Raspberry Pi 4B control board, an RGB camera (resolution 640×480), a 4-channel line-tracking sensor, an LFD-01 PMM servo, and a Mecanum wheel drive system. The trained policy network is deployed on the Raspberry Pi 4B, and control commands are inferred in real-time using Python.
Expert Data: The vehicle is guided to autonomously drive along the track centerline 10 times using the 4-channel line-tracking sensor, collecting continuous state trajectories (including camera images) as expert state inputs.
Experimental Environment: As shown in Fig. 5, the track is 200cm $\times$ 150 cm, including straight and curved sections.
Experimental Results: Fig. 5 illustrates the actual experimental scenario layout and the vehicle’s driving states at multiple positions under RFRLF control. The experimental results demonstrate that, in the absence of reward signals and expert actions, RFRLF can optimize the policy by predicting the error between the state and the expert state, enabling the vehicle to drive stably along the track centerline with an average lateral error of 3.2 cm (standard deviation ±1.5 cm).
Figure 5: The central area is a 200×150 cm autonomous driving track, with the starting point and endpoint marked. Surrounding the track are top-down views of the vehicle driving at different positions, corresponding to the blue marking points on the track.
IV-E Performance Evaluation of Target State Prediction Network
@@
The key reason why the TSPN can achieve high-precision predictions lies in the action injection layer in its design. By injecting action information into the state features, the network is able to capture the impact of actions on state changes. In addition, the feature reconstruction and feature extraction modules not only effectively preserve the spatial resolution of the features, but also enable the network to capture richer contextual information, thereby further improving the prediction accuracy.
@@
IV-F Parameter Sensitivity Analysis
In the Carla environment, we implement 5 comparison schemes by setting various hyperparameters and making changes to components. In Table II, “Partial Freeze (PF)” indicates freezing some layers (input embedding layer and feature extraction layer) of the TSPN. This is to analyze the impact of free adjustment of other layers on performance while retaining some feature extraction capabilities. In order to analyze the impact of temperature parameter “$T$=1.0” and scheduled sampling rate “$\epsilon$=0.5” on the RFRLF method, we also set $T$ in {0.5,2.0} and $\epsilon$ in {0.3,0.7} , where $T$ is the temperature parameter during data collection and $\epsilon$ is the scheduled sampling rate after the TSPN training epoch is greater than 2. These parameters are selected to observe the impact of smaller and larger temperatures and sampling rates on model exploration and training stability.
@@ IV-F Parameter Sensitivity Analysis
Overall, the performance of RFRLF is comparable to the baseline with altered hyperparameters and outperforms the baseline with partial component structure changes in terms of Mean Value and IQM. Specifically, by comparing with the “PF” baseline, it is found that after freezing part of the TSPN, the network performance does not change much. The frozen input embedding and feature extraction layers preserve the network’s feature extraction capabilities, while the unfrozen components remain adaptable during training, collectively ensuring the effectiveness of the method.
For the temperature parameter, the “$T$=2.0” baseline performs best. This is because when training the TSPN, a higher temperature coefficient enhances exploration, enabling the agent to explore more states, thereby providing the TSPN with richer data. However, as the temperature coefficient increases, instability also increases, which means that the $T$ should not be too high, and a balance needs to be found between exploration and stability. Experimental results show that, “$T$=1.0” as the default value is a suitable compromise. In addition, by comparing the “$T$=0.5”, “$T$=2.0”, “$\epsilon$=0.3”, and “$\epsilon$=0.7” baselines, we find that hyperparameters have a certain impact on the performance of RFRLF, but in general, RFRLF is not very sensitive to changes in hyperparameters. This shows that RFRLF can adapt to different hyperparameter settings.
@@ V Conclusion
This paper proposes a novel Reward-Free Reinforcement Learning Framework (RFRLF) that optimizes decision-making processes without requiring manually designed reward signals or explicit action information, through a dual-module architecture comprising the Target State Prediction Network (TSPN) and the Reward-Free State-Guided Policy Network (RFSGPN). The TSPN predicts the next state based on the agent’s current state and executed action, while the RFSGPN leverages the prediction results and expert demonstration data (without action information) for policy learning. Experimental results demonstrate that RFRLF outperforms multiple reward-based and reward-free reinforcement learning baselines in Carla and Autocar simulation environments, achieving significant advantages in maximum, average, and IQM metrics. Furthermore, in real-world validation on autonomous driving tracks, the framework successfully achieves path-following tasks without manually designed reward signals, demonstrating its effectiveness in practical scenarios.
Although RFRLF demonstrates strong performance in both simulated and real-world environments, its effectiveness in complex scenarios (e.g., extreme lighting conditions, dynamic obstacles) requires further improvement. Future research will explore state prediction enhancement methods based on multi-modal perception and extend the framework to multi-agent policy optimization scenarios, thereby improving performance in complex environments.